What is the answer to 8m-5=5m+7
1
Expert's answer
2013-01-16T08:31:52-0500
8m-5=5m+7
Let's move all terms containing m to the left part, all the rest to the right. Keep in mind that if you move a term from one part of equation to another you should change the sign of it:
8m-5m=7+5
3m=12
Now we need to divide both parts by 3, we obtain:
m=4